# On Random Fields Associated with Analytic Wavelet Transform

### v1

@@ 5.3. Proof of Corollary 1
$p\left(r_{1:n},\theta_{1:n};|\mathbf{W}_{\Phi}|,\Theta\right)$
(56) $=$ $\frac{r_{1}r_{2}\cdots r_{n}}{\pi^{n}\textup{det}(\Gamma)}\textup{exp}\left(-\begin{bmatrix}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\\ \vdots\\ r_{n}e^{i\theta_{n}}\end{bmatrix}^{*}\Gamma^{-1}\left[\begin{array}[]{cc}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\\ \vdots\\ r_{n}e^{i\theta_{n}}\end{array}\right]\right),$
@@
$\begin{bmatrix}e^{-i\theta_{1}}&e^{-i\theta_{2}}\end{bmatrix}\mathbf{M}\begin{bmatrix}e^{i\theta_{1}}\\ e^{i\theta_{2}}\end{bmatrix}$
$=$ $M_{11}+M_{22}+\left(M_{12}e^{i(\theta_{2}-\theta_{1})}+M_{21}e^{-i(\theta_{2}-\theta_{1})}\right)$
@@
(57) $J(\mathbf{M})=$ $e^{-(M_{11}+M_{22})}$
$\times\int_{0}^{2\pi}\int_{0}^{2\pi}\exp\left(-2\Re\left(M_{12}e^{i(\theta_{2}-\theta_{1})}\right)\right)d\theta_{1}d\theta_{2}.$
@@
$J(\mathbf{M})=$ $e^{-(M_{11}+M_{22})}$
$\times\int_{0}^{2\pi}\hskip-5.69046pt\int_{0}^{2\pi}\hskip-5.69046pt\exp\left(-2|M_{12}|\cos(\theta_{2}-\theta_{1}+\phi)\right)d\theta_{1}d\theta_{2}.$
@@
$\mathbf{M}=$ $\textup{diag}(r_{1:2})\Gamma^{-1}\textup{diag}(r_{1:2})$
$=$ $\frac{1}{\textup{det}(\Gamma)}\begin{bmatrix}r_{1}^{2}\Gamma_{22}&-r_{1}r_{2}\Gamma_{12}\\ -r_{1}r_{2}\overline{\Gamma_{12}}&r_{2}^{2}\Gamma_{11}\end{bmatrix}$,
@@
$p\left(r_{1:2};|\mathbf{W}_{\Phi}|\right)$
$=$ $\frac{4r_{1}r_{2}}{\textup{det}(\Gamma)}\exp\left(-\frac{r_{1}^{2}\Gamma_{22}+r_{2}^{2}\Gamma_{11}}{\textup{det}(\Gamma)}\right)I_{0}\left(\frac{2|\Gamma_{12}|}{\textup{det}(\Gamma)}r_{1}r_{2}\right).$
@@ 5.4. Proof of Proposition 2
$N_{R}=$ $\Re\left(e^{-i\varphi}W_{\Phi}(t,s)\right)$
(60) $=$ $\cos(\varphi)\Re(W_{\Phi}(t,s))+\sin(\varphi)\Im(W_{\Phi}(t,s))$
@@
$N_{I}=$ $\Im\left(e^{-i\varphi}W_{\Phi}(t,s)\right)$
(61) $=$ $\cos(\varphi)\Im(W_{\Phi}(t,s))-\sin(\varphi)\Re(W_{\Phi}(t,s)).$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)$
$=$ $\mathbb{P}\left(\left|\frac{2N_{R}}{|W_{f}(t,s)|}+\frac{N^{2}_{R}+N^{2}_{I}}{|W_{f}(t,s)|^{2}}\right|>\varepsilon\right)$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)\leq$ $\textup{exp}\left(-\frac{\varepsilon^{2}}{32}\frac{|W_{f}(t,s)|^{2}}{\sigma_{s}^{2}}\right)$
(65) $+$ $\textup{exp}\left(-\frac{\varepsilon}{4}\frac{|W_{f}(t,s)|^{2}}{\sigma_{s}^{2}}\right).$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)$
$=$ $\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}-1\right|>\left(\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}+1\right)^{-1}\varepsilon\right)$
@@ 5.6. Proof of Proposition 4
Sard’s Theorem [32, Chapter 6] Let $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a function of class $C^{k}$, where $k\geq\max(n-m+1,1)$. Then the set of critical values of $g$, that is,
@@
Regular Level Set Theorem [33, Corollary A.26] Let $M$ and $N$ be smooth manifolds, and let $\Phi:M\rightarrow N$ be a smooth map. Every regular level set of $\Phi$ is a properly embedded submanifold of $M$ whose codimension is equal to $dim(N)$.
@@ 5.7. Proof of Proposition 5
$p(\theta;\Theta_{Y}(t,s))$
$=$ $\frac{1}{\pi\sigma_{s}^{2}}\int_{0}^{\infty}\textup{exp}\left\{-\frac{\left(r\cos\theta-\mathfrak{R}(W_{f}(t,s))\right)^{2}}{\sigma_{s}^{2}}\right$.
$\hskip 62.59596pt-\left.\frac{\left(r\sin\theta-\mathfrak{I}(W_{f}(t,s))\right)^{2}}{\sigma_{s}^{2}}\right\}rdr$
@@
$\times\int_{0}^{\infty}\textup{exp}\left\{\frac{-r^{2}+2r\mathfrak{R}(W_{f}(t,s))\cos\theta}{\sigma_{s}^{2}}\right$.
(71) $\hskip 73.97733pt\left.+\frac{2r\mathfrak{I}(W_{f}(t,s))\sin\theta}{\sigma_{s}^{2}}\right\}rdr$
@@
(72) $p(\theta;\Theta_{Y}(t,s))$
@@ 5.8. Proof of Corollary 2
$p\left(\theta_{1:n};\Theta_{\Phi}\right)=$ $\frac{1}{\pi^{n}\textup{det}(\Gamma)}\int_{0}^{\infty}\int_{S_{+}^{n-1}}\rho^{n}\left(\overset{n}{\underset{\ell=1}{\prod}}\omega_{\ell}\right)$
$\times\exp\left(-\rho^{2}\omega^{\top}\mathbf{M}\omega\right)\rho^{n-1}d\omega d\rho,$
@@
$2\pi^{2}\textup{det}(\Gamma)p\left(\theta_{1:2};\Theta_{\Phi}\right)$
$=$ $\int_{0}^{\pi/2}\hskip-5.69046pt\frac{\sin u\cos u}{\left(M_{22}\sin^{2}u+2\Re(M_{12})\cos u\sin u+M_{11}\cos^{2}u\right)^{2}}du$
@@
$2\pi^{2}\textup{det}(\Gamma)p\left(\theta_{1:2};\Theta_{\Phi}\right)$
$=$ $\int_{0}^{\infty}\frac{v}{\left(M_{22}v^{2}+2\Re(M_{12})v+M_{11}\right)^{2}}dv$
$=$ $\frac{1}{2(M_{11}M_{22}-|\Re(M_{12})|^{2})}\Big{.}$
$\Big{.}-\frac{\Re(M_{12})}{2(M_{11}M_{22}-|\Re(M_{12})|^{2})^{3/2}}$
@@
$p\left(\theta_{1:2};\Theta_{\Phi}\right)$
$=$ $\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)}{\Gamma_{11}\Gamma_{22}-|\Re(\Gamma_{12}e^{i(\theta_{2}-\theta_{1})})|^{2}}$
@@ 5.9. Proof of Proposition 6
$|W_{Y}(t,s)|^{2}$
$=$ $\left(\mathfrak{R}(W_{f}(t,s))\right)^{2}+2\mathfrak{R}(W_{f}(t,s))\mathfrak{R}(W_{\Phi}(t,s))$
@@
$\left(\Box(W_{\Phi}(t,s))\right)^{2}=s\int_{\mathbb{R}}|\widehat{\Box(\psi)}(su)|^{2}$
$+s\int_{\mathbb{R}^{2}}^{\prime}e^{it(u+v)}\overline{\widehat{\Box(\psi)}}(su)\overline{\widehat{\Box(\psi)}}(sv)Z_{F}(du)Z_{F}(dv)F(du),$
@@
$|W_{Y}(t,s)|^{2}-\mathbb{E}\left[|W_{Y}(t,s)|^{2}\right]$
$=$ $2\Re(W_{f}(t,s))\sqrt{s}\int_{\mathbb{R}}e^{it\lambda}\overline{\widehat{\Re(\psi)}}(s\lambda)Z_{F}(d\lambda)$
@@
$=$ $4\underset{\Box,\triangle\in\{\Re,\Im\}}{\sum}\Box(W_{f}(t_{1},s_{1}))\triangle(W_{f}(t_{2},s_{2}))$
$\times\sqrt{s_{1}s_{2}}\int_{\mathbb{R}}e^{i(t_{1}-t_{2})\lambda}\overline{\widehat{\Box(\psi)}}(s_{1}\lambda)\widehat{\triangle(\psi)}(s_{2}\lambda)F(d\lambda)$
@@ 5.10. Proof of Proposition 7
(80) $=$ $\frac{1}{\pi^{2}\textup{det}(\Gamma)|w_{1}|^{2}}\textup{exp}\left(-\left(\left[\begin{array}[]{cc}w_{1}\\ z/w_{1}\end{array}\right]-\mathbf{W}_{f}\right)^{*}\right$.
(83) $\left.\Gamma^{-1}\left(\left[\begin{array}[]{cc}w_{1}\\ z/w_{1}\end{array}\right]-\mathbf{W}_{f}\right)\right),\ w_{1},z\in\mathbb{C},$
which implies that the marginal probability density function of $W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})$ has the integral form
@@
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}p\left(re^{i\theta},z;W_{Y}(t_{1},s_{1}),W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)$
$rd\theta dr$
(86) $=$ $\int_{0}^{\infty}\int_{0}^{2\pi}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\textup{exp}\left(-\left(\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]-\mathbf{W}_{f}\right)^{*}\right$.
(89) $\left.\Gamma^{-1}\left(\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]-\mathbf{W}_{f}\right)\right)d\theta dr.$
@@ 5.11. Proof of Corollary 4
$p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)$
[…]
(95) $\left.\left.\Gamma^{-1}\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]\right)d\theta\right\}dr,$
@@
(100) $\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]$
(101) $=$ $\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}-2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}.$
@@
(102) $p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)$
$=$ $\left\{\int_{0}^{2\pi}\textup{exp}\left(\frac{2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}\right)d\theta\right\}$
@@
$\int_{0}^{2\pi}\textup{exp}\left(\frac{2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}\right)d\theta$
$=$ $\int_{0}^{2\pi}\textup{exp}\left(\frac{2|z\Gamma_{12}|}{\det(\Gamma)}\mathfrak{R}\left(e^{i(\nu-2\theta)}\right)\right)d\theta$
@@
$\int_{0}^{\infty}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr$
(104) $=$ $\frac{1}{\pi^{2}\textup{det}(\Gamma)}K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right).$
@@
$p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)$
(105) $=$ $\frac{2}{\pi\textup{det}(\Gamma)}I_{0}\left(\frac{2|\Gamma_{12}|}{\det(\Gamma)}|z|\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right)$
@@
$\mathbb{E}\left[\left|W_{\Phi}(t_{1},s_{1})\right|\left|W_{\Phi}(t_{2},s_{2})\right|\right]$
$=$ $\int_{0}^{2\pi}\int_{0}^{\infty}r^{2}p\left(r;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)drd\theta$

### v2

@@ 5.3. Proof of Corollary 1
(61) $p\left(r_{1:n},\theta_{1:n};|\mathbf{W}_{\Phi}|,\Theta\right)=\frac{r_{1}r_{2}\cdots r_{n}}{\pi^{n}\textup{det}(\Gamma)}\textup{exp}\left(-\begin{bmatrix}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\\ \vdots\\ r_{n}e^{i\theta_{n}}\end{bmatrix}^{*}\Gamma^{-1}\left[\begin{array}[]{cc}r_{1}e^{i\theta_{1}}\\ r_{2}e^{i\theta_{2}}\\ \vdots\\ r_{n}e^{i\theta_{n}}\end{array}\right]\right),$
@@
$\begin{bmatrix}e^{-i\theta_{1}}&e^{-i\theta_{2}}\end{bmatrix}\mathbf{M}\begin{bmatrix}e^{i\theta_{1}}\\ e^{i\theta_{2}}\end{bmatrix}$= $M_{11}+M_{22}+\left(M_{12}e^{i(\theta_{2}-\theta_{1})}+M_{21}e^{-i(\theta_{2}-\theta_{1})}\right)$
@@
(62) $J(\mathbf{M})=$ $e^{-(M_{11}+M_{22})}\int_{0}^{2\pi}\int_{0}^{2\pi}\exp\left(-2\Re\left(M_{12}e^{i(\theta_{2}-\theta_{1})}\right)\right)d\theta_{1}d\theta_{2}.$
@@
$J(\mathbf{M})=$ $e^{-(M_{11}+M_{22})}\int_{0}^{2\pi}\hskip-5.69046pt\int_{0}^{2\pi}\hskip-5.69046pt\exp\left(-2|M_{12}|\cos(\theta_{2}-\theta_{1}+\phi)\right)d\theta_{1}d\theta_{2}.$
@@
$\mathbf{M}=\textup{diag}(r_{1:2})\Gamma^{-1}\textup{diag}(r_{1:2})=\frac{1}{\textup{det}(\Gamma)}\begin{bmatrix}r_{1}^{2}\Gamma_{22}&-r_{1}r_{2}\Gamma_{12}\\ -r_{1}r_{2}\overline{\Gamma_{12}}&r_{2}^{2}\Gamma_{11}\end{bmatrix}$,
@@
$p\left(r_{1:2};|\mathbf{W}_{\Phi}|\right)=\frac{4r_{1}r_{2}}{\textup{det}(\Gamma)}\exp\left(-\frac{r_{1}^{2}\Gamma_{22}+r_{2}^{2}\Gamma_{11}}{\textup{det}(\Gamma)}\right)I_{0}\left(\frac{2|\Gamma_{12}|}{\textup{det}(\Gamma)}r_{1}r_{2}\right).$
@@ 5.4. Proof of Proposition 2
(65) $N_{R}=\Re\left(e^{-i\varphi}W_{\Phi}(t,s)\right)=\cos(\varphi)\Re(W_{\Phi}(t,s))+\sin(\varphi)\Im(W_{\Phi}(t,s))$
@@
(66) $N_{I}=$ $\Im\left(e^{-i\varphi}W_{\Phi}(t,s)\right)=\cos(\varphi)\Im(W_{\Phi}(t,s))-\sin(\varphi)\Re(W_{\Phi}(t,s)).$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)=$ $\mathbb{P}\left(\left|\frac{2N_{R}}{|W_{f}(t,s)|}+\frac{N^{2}_{R}+N^{2}_{I}}{|W_{f}(t,s)|^{2}}\right|>\varepsilon\right)$
@@
(70) $\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)\leq\textup{exp}\left(-\frac{\varepsilon^{2}}{32}\frac{|W_{f}(t,s)|^{2}}{\sigma_{s}^{2}}\right)+\textup{exp}\left(-\frac{\varepsilon}{4}\frac{|W_{f}(t,s)|^{2}}{\sigma_{s}^{2}}\right).$
@@
$\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|^{2}}{|W_{f}(t,s)|^{2}}-1\right|>\varepsilon\right)=$ $\mathbb{P}\left(\left|\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}-1\right|>\left(\frac{|W_{Y}(t,s)|}{|W_{f}(t,s)|}+1\right)^{-1}\varepsilon\right)$
@@ 5.6. Proof of Proposition 4
Sard’s Theorem [37, Chapter 6] Let $g:\mathbb{R}^{n}\to\mathbb{R}^{m}$ be a function of class $C^{k}$, where $k\geq\max(n-m+1,1)$. Then the set of critical values of $g$, that is,
@@
Regular Level Set Theorem [38, Corollary A.26] Let $M$ and $N$ be smooth manifolds, and let $\Phi:M\rightarrow N$ be a smooth map. Every regular level set of $\Phi$ is a properly embedded submanifold of $M$ whose codimension is equal to $dim(N)$.
@@ 5.7. Proof of Proposition 5
(76) $p(\theta;\Theta_{Y}(t,s))$
$=$ $\frac{1}{\pi\sigma_{s}^{2}}\int_{0}^{\infty}\textup{exp}\left\{-\frac{\left(r\cos\theta-\mathfrak{R}(W_{f}(t,s))\right)^{2}}{\sigma_{s}^{2}}-\frac{\left(r\sin\theta-\mathfrak{I}(W_{f}(t,s))\right)^{2}}{\sigma_{s}^{2}}\right\}rdr$
@@
$\times\int_{0}^{\infty}\textup{exp}\left\{\frac{-r^{2}+2r\mathfrak{R}(W_{f}(t,s))\cos\theta}{\sigma_{s}^{2}}+\frac{2r\mathfrak{I}(W_{f}(t,s))\sin\theta}{\sigma_{s}^{2}}\right\}rdr$
@@
(77) $p(\theta;\Theta_{Y}(t,s))=$ $\frac{1}{\pi\sigma_{s}^{2}}\textup{exp}\left\{-\frac{1}{\sigma_{s}^{2}}\left|W_{f}(t,s)\right|^{2}\right\}$
@@ 5.8. Proof of Corollary 2
$p\left(\theta_{1:n};\Theta_{\Phi}\right)=\frac{1}{\pi^{n}\textup{det}(\Gamma)}\int_{0}^{\infty}\int_{S_{+}^{n-1}}\rho^{n}\left(\overset{n}{\underset{\ell=1}{\prod}}\omega_{\ell}\right)\exp\left(-\rho^{2}\omega^{\top}\mathbf{M}\omega\right)\rho^{n-1}d\omega d\rho,$
@@
$2\pi^{2}\textup{det}(\Gamma)p\left(\theta_{1:2};\Theta_{\Phi}\right)=$ $\int_{0}^{\pi/2}\hskip-5.69046pt\frac{\sin u\cos u}{\left(M_{22}\sin^{2}u+2\Re(M_{12})\cos u\sin u+M_{11}\cos^{2}u\right)^{2}}du$
@@
$2\pi^{2}\textup{det}(\Gamma)p\left(\theta_{1:2};\Theta_{\Phi}\right)=$ $\int_{0}^{\infty}\frac{v}{\left(M_{22}v^{2}+2\Re(M_{12})v+M_{11}\right)^{2}}dv$
$=$ $\frac{1}{2(M_{11}M_{22}-|\Re(M_{12})|^{2})}-\frac{\Re(M_{12})}{2(M_{11}M_{22}-|\Re(M_{12})|^{2})^{3/2}}$
@@
$p\left(\theta_{1:2};\Theta_{\Phi}\right)=$ $\frac{1}{4\pi^{2}}\frac{\textup{det}(\Gamma)}{\Gamma_{11}\Gamma_{22}-|\Re(\Gamma_{12}e^{i(\theta_{2}-\theta_{1})})|^{2}}$
@@ 5.9. Proof of Proposition 6
$|W_{Y}(t,s)|^{2}=$ $\left(\mathfrak{R}(W_{f}(t,s))\right)^{2}+2\mathfrak{R}(W_{f}(t,s))\mathfrak{R}(W_{\Phi}(t,s))$
@@
$\left(\Box(W_{\Phi}(t,s))\right)^{2}=s\int_{\mathbb{R}}|\widehat{\Box(\psi)}(su)|^{2}+s\int_{\mathbb{R}^{2}}^{\prime}e^{it(u+v)}\overline{\widehat{\Box(\psi)}}(su)\overline{\widehat{\Box(\psi)}}(sv)Z_{F}(du)Z_{F}(dv)F(du),$
@@
$|W_{Y}(t,s)|^{2}-\mathbb{E}\left[|W_{Y}(t,s)|^{2}\right]=$ $2\Re(W_{f}(t,s))\sqrt{s}\int_{\mathbb{R}}e^{it\lambda}\overline{\widehat{\Re(\psi)}}(s\lambda)Z_{F}(d\lambda)$
@@
$=$ $4\underset{\Box,\triangle\in\{\Re,\Im\}}{\sum}\Box(W_{f}(t_{1},s_{1}))\triangle(W_{f}(t_{2},s_{2}))\sqrt{s_{1}s_{2}}\int_{\mathbb{R}}e^{i(t_{1}-t_{2})\lambda}\overline{\widehat{\Box(\psi)}}(s_{1}\lambda)\widehat{\triangle(\psi)}(s_{2}\lambda)F(d\lambda)$
@@ 5.10. Proof of Proposition 7
(87) $=$ $\frac{1}{\pi^{2}\textup{det}(\Gamma)|w_{1}|^{2}}\textup{exp}\left(-\left(\left[\begin{array}[]{cc}w_{1}\\ z/w_{1}\end{array}\right]-\mathbf{W}_{f}\right)^{*}\Gamma^{-1}\left(\left[\begin{array}[]{cc}w_{1}\\ z/w_{1}\end{array}\right]-\mathbf{W}_{f}\right)\right)$
for any $w_{1},z\in\mathbb{C}$, which implies that the marginal probability density function of $W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})$ has the integral form
@@
$=$ $\int_{0}^{\infty}\int_{0}^{2\pi}p\left(re^{i\theta},z;W_{Y}(t_{1},s_{1}),W_{Y}(t_{1},s_{1})W_{Y}(t_{2},s_{2})\right)rd\theta dr$
(92) $=$ $\int_{0}^{\infty}\int_{0}^{2\pi}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\textup{exp}\left(-\left(\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]-\mathbf{W}_{f}\right)^{*}\Gamma^{-1}\left(\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]-\mathbf{W}_{f}\right)\right)d\theta dr.$
@@ 5.11. Proof of Corollary 4
(93) $p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)$
(98) $=$ $\int_{0}^{\infty}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\left\{\int_{0}^{2\pi}\textup{exp}\left(-\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]\right)d\theta\right\}dr,$
@@
(103) $\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]^{*}\Gamma^{-1}\left[\begin{array}[]{cc}re^{i\theta}\\ zr^{-1}e^{-i\theta}\end{array}\right]=\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}-2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}.$
@@
(104) $p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)=$ $\left\{\int_{0}^{2\pi}\textup{exp}\left(\frac{2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}\right)d\theta\right\}$
@@
$\int_{0}^{2\pi}\textup{exp}\left(\frac{2\mathfrak{R}\left(z\Gamma_{12}e^{-i2\theta}\right)}{\det(\Gamma)}\right)d\theta=$ $\int_{0}^{2\pi}\textup{exp}\left(\frac{2|z\Gamma_{12}|}{\det(\Gamma)}\mathfrak{R}\left(e^{i(\nu-2\theta)}\right)\right)d\theta$
@@
(106) $\int_{0}^{\infty}\frac{1}{\pi^{2}\textup{det}(\Gamma)r}\textup{exp}\left(-\frac{\Gamma_{22}r^{2}+\Gamma_{11}|z|^{2}r^{-2}}{\det(\Gamma)}\right)dr=\frac{1}{\pi^{2}\textup{det}(\Gamma)}K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right).$
@@
(107) $p\left(z;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)=\frac{2}{\pi\textup{det}(\Gamma)}I_{0}\left(\frac{2|\Gamma_{12}|}{\det(\Gamma)}|z|\right)K_{0}\left(\frac{2\sqrt{\Gamma_{11}\Gamma_{22}}}{\det(\Gamma)}|z|\right)$
@@
$\mathbb{E}\left[\left|W_{\Phi}(t_{1},s_{1})\right|\left|W_{\Phi}(t_{2},s_{2})\right|\right]=$ $\int_{0}^{2\pi}\int_{0}^{\infty}r^{2}p\left(r;W_{\Phi}(t_{1},s_{1})W_{\Phi}(t_{2},s_{2})\right)drd\theta$